If a culture has an instantaneous growth rate constant of 0.03235129 min , what is the doubling time of the culture? Express your answer in minutes using exponential notation. o 48.2 minutes would be 4.82e1 . Round vour answer to two decimal places o 32.67 minutes would be 2.27e1
Answer #1

Answer:

Exponential growth equation: eμtg = 2---(1) where,

μ = instantaneous growth-rate constant

tg = Doubling time

Given that,

μ = 0.03235129 min-1

Using the Eqtn(1), ln(2) = ln(eμtg)

\Rightarrow ln (2) = μtg =(0.03235129 min-1) (tg)

\Rightarrow tg = ln (2) / (0.03235129 min-1) = 0.693 / 0.03235129 = 21.42 minutes = 2.14e1